# Arithmetic of cuts in ordered abelian groups and of ideals over valuation rings

### v2

@@ 2.3. Subtraction of final segments
The final segment $S_{2}-_{\rm mc}S_{1}$ is closed, except in the following cases. If $S_{2}=\alpha^{+}$ and $S_{1}=\beta^{-}$ for some $\alpha,\beta\in G$ and $\alpha$ has no immediate successor in $G$, then $\alpha+\beta$ has no immediate successor in $G$ and therefore, $S_{2}-_{\rm mc}S_{1}=(\alpha+\beta)^{+}$ is not closed. If $S_{2}$ and $S_{1}$ have no infimum in $G$, then $S_{2}-_{\rm mc}S_{1}$ may or may not have an infimum in $G$. For instance, if $G=\mathbb{Q}$ and $S_{1}=S_{2}=\{\alpha\in\mathbb{Q}\mid\alpha>\pi\}$, then $\Delta S_{1}=\{\alpha\in\mathbb{Q}\mid\alpha>-\pi\}$ and $S_{2}-_{\rm mc}S_{1}=0^{+}$ which is not closed although $S_{1}$ and $S_{2}$ are.
The question arises whether the difference operation $-_{\rm mc}$ provides some positive answer to our questions (QFS1) and (QFS2). At this point, we can give a complete answer to these questions in the case where $S_{1}$ is principal:
Lemma 2.12.
Take two final segments $S_{1}$ and $S_{2}$ of $G$. If $S_{1}$ is principal, say $S_{1}=\alpha^{-}$, then
[…]
and $T=S_{2}\,-_{\rm mc}\,S_{1}=S_{2}-\alpha$ is the unique solution of $S_{1}+T=S_{2}\,$. Hence
$S_{1}\,-_{\rm fs}\,S_{2}\>=\>S_{2}-\alpha\>.$
Proof.
For $S_{1}=\alpha^{-}$, the first equation follows from equation (13). Now equation (16) implies that $S_{1}+(S_{2}-_{\rm mc}S_{1})=S_{2}\,$. The uniqueness of $T$ is stated in part 2) of Proposition 2.5. ∎
It now remains to deal with the case where $S_{1}$ is nonprincipal. This case is much more complex and we need some preparation.
@@ 2.4. Invariance groups
Take any non-empty subset $M\subseteq G$. Then we have:
1)  ${\mathcal{G}}(M)$ is a subgroup of $G$.
@@
4)  If $0\in M$, then ${\mathcal{G}}(M)\subseteq M$.
@@
(19) ${\mathcal{G}}(\alpha^{-})\>=\>{\mathcal{G}}(\alpha^{+})\>=\>{\mathcal{G}}(\{\alpha\})\>=\>\{0\}\>,$
which shows that the invariance group of every final segment that has an infimum in $G$ is trivial.
@@
4):  If $0\in M$, then ${\mathcal{G}}(M)=0+{\mathcal{G}}(M)\subseteq M+{\mathcal{G}}(M)=M$.
@@
6):  Assume that $\alpha\in G$ is the infimum of $M$ in $G$. Then for every positive $\gamma\in G$ there is $\beta\in M$ such that $0\leq\beta-\alpha<\gamma$. Hence $\beta-\\ gammaamma<\alpha$ so that $\beta-\gamma\notin M$ and thus $\gamma\notin{\mathcal{G}}(M)$ by (18). This proves that ${\mathcal{G}}(M)=\{0\}$. ∎
@@
for every two convex subgroups $H_{1}$ and $H_{2}\,$.
@@
${\mathcal{G}}(S_{1}+S_{2})\>=\>{\mathcal{G}}(S_{1})\cup{\mathcal{G}}(S_{2})\>.$
3)  If $\alpha\in G$ and $S$ is a final segment of $G$, then $\alpha+S=\alpha^{-}+S$ is again a final segment of $G$, and
@@
3):  The proof of the first statement is straightforward. It follows that ${\mathcal{G}}(\alpha+S)={\mathcal{G}}(\alpha^{-}+S)={\mathcal{G}}(\alpha^{-})\cup{\mathcal{G}}(S)=\{0\}\cup{\mathcal{G}}(S)={\mathcal{G}}(S)$, where we have used part 2) of our lemma and equation (19) of Lemma 2.13.
@@
6):  We use that ${\mathcal{G}}(S)^{-}$ is the disjoint union of ${\mathcal{G}}(S)$ and ${\mathcal{G}}(S)^{+}$. If $\alpha\in{\mathcal{G}}(S)$, then $\alpha+S=S$. If $\alpha\in{\mathcal{G}}(S)^{+}$, then $\alpha>0$ and therefore, $\alpha+S\subseteq S$; but as $\alpha\notin{\mathcal{G}}(S)$, $\alpha+S\neq S$. This proves the implications $\Leftarrow$” in both statements.
@@
Note that the set $\{na\mid\alpha\in S\}$ is a final segment if and only if $G$ is $n$-divisible; in general, $nS$ is the smallest final segment that contains $\{n\alpha\mid\alpha\in S\}$. By a repeated application of part 2) of the last lemma, we obtain:
@@
From part 1) of Lemma 2.14 we obtain:
@@
Take final segments $S_{1}\,$ $S_{2}$ and $T$ such that $S_{1}+T=S_{2}$. Then ${\mathcal{G}}(S_{1})\subseteq{\mathcal{G}}(S_{2})$, and if ${\mathcal{G}}(S_{1})\subsetneq{\mathcal{G}}(S_{2})$, then ${\mathcal{G}}(T)={\mathcal{G}}(S_{2})$.
@@
Take final segments $S$, $S_{1}$ and $S_{2}$ of $G$.
@@
2) We have that
${\mathcal{G}}(S_{2}\,-_{\rm mc}\,S_{1})\>=\>{\mathcal{G}}(S_{1})+{\mathcal{G}}(S_{2})\>.$
3) If ${\mathcal{G}}(S)\neq\{0\}$ then $S$ is nonprincipal.
@@
1): The first and second equality are special cases of part 3) of Lemma 2.13. The equality ${\mathcal{G}}(S_{1})={\mathcal{G}}(\Delta S_{1})$ is seen as follows. If $S_{1}$ is nonprincipal, then by part 2) of Lemma 2.10, $\Delta S_{1}=-S_{1}^{c}$ and the equality follows from what we have already proved. If $S_{1}$ is principal, say $S_{1}=\alpha^{-}$ for some $\alpha\in G$, then $\Delta S_{1}=(-\alpha)^{-}$ by (14), whence ${\mathcal{G}}(\Delta S_{1})=\{0\}={\mathcal{G}}(S_{1})$ by (19).
2): This follows from part 1) of our lemma together with parts 1) and 4) of Lemma 2.10 and part 2) of Lemma 2.14.
3): If ${\mathcal{G}}(S)\neq\{0\}$, then there is a negative element $\alpha\in{\mathcal{G}}(S)$ so that for every $\beta\in S\,$, $\beta>\alpha+\beta\in S\,$; hence $S$ has no smallest element. ∎
In order to give answers to questions (QFS1), (QFS2) and (QFS3) in the case where ${\mathcal{G}}(S_{1})={\mathcal{G}}(S_{2})=\{0\}$ in Section 2.5, we need some more preparation. We set $G^{>0}:=\{\gamma\in G\mid\gamma>0\}=0^{+}$.
Lemma 2.18.
Proof.
If $\gamma=\sigma-\sigma^{\prime}$ with $\sigma\in S$ and $\sigma^{\prime}\in S^{c}$, then $\sigma-\gamma=\sigma^{\prime}\notin S$, so $-\gamma\notin{\mathcal{G}}(S)$. Since ${\mathcal{G}}(S)$ is a subgroup of $G$, this implies that $\gamma\notin{\mathcal{G}}(S)$. Conversely, if $H$ is a subgroup of $G$ properly containing ${\mathcal{G}}(S)$, then there is $\gamma\in H^{>0}\setminus{\mathcal{G}}(S)$. Then $-\gamma\notin{\mathcal{G}}(S)$ and there is $\sigma\in S$ such that $\sigma-\gamma\notin S$. This implies $\sigma^{\prime}=\sigma-\gamma\in S^{c}$, whence $\gamma=\sigma-\sigma^{\prime}\in S-S^{c}$.
Now we prove equation (20). Since $\alpha<\beta$ whenever $\alpha\in S^{c}$ and $\beta\in S$, it follows that $S-S^{c}\subseteq G^{>0}\setminus{\mathcal{G}}(S)$. For the converse, assume that $\gamma\in G^{>0}\setminus{\mathcal{G}}(S)$. Then there is some $\alpha\in S^{c}$ such that $\alpha+\gamma\in S$ and therefore, $\gamma\in S-S^{c}$. ∎
Corollary 2.19.
1) We have
$S_{1}\,-_{\rm mc}\,S_{1}\>=\>S_{1}-S_{1}^{c}\>=\>{\mathcal{G}}(S_{1})^{+}\>,$
so that
(21) $S_{1}\,+\,(S_{2}\,-_{\rm mc}\,S_{1})\>=\>{\mathcal{G}}(S_{1})^{+}+S_{2}\>\subseteq S_{2}\>.$
Proof.
1): This follows from equation (15) in part 2) of Lemma 2.10 together with Lemma 2.18. The inequality holds since ${\mathcal{G}}(S_{1})^{+}$ only contains positive elements.
2): This follows from part 1) together with part 3) of Lemma 2.6
3): Since $S_{1}$ is nonprincipal, we can apply part 1). As ${\mathcal{G}}(S_{1})\subsetneq{\mathcal{G}}(S_{2})$, there is some $\alpha\in{\mathcal{G}}(S_{1})^{+}\cap{\mathcal{G}}(S_{2})$. We obtain that $S_{2}=\alpha+S_{2}\subseteq{\mathcal{G}}(S_{1})^{+}+S_{2}\,$. In view of (21), this proves statement 3). ∎
Finally, for application in Section 3.3 we prove the following facts:
Proposition 2.20.
Take final segments $S$, $S_{1}$ and $S_{2}$ of $G$.
1) We have that ${\mathcal{G}}(S)^{+}$ does not contain $S$ if and only if $0\in S$.
Proof.
1): If ${\mathcal{G}}(S)^{+}$ does not contain $\alpha+S$, then $S\setminus{\mathcal{G}}(S)^{+}\neq\emptyset$, so there is some $\alpha\in S\cap{\mathcal{G}}(S)$. Since ${\mathcal{G}}(S)$ is a group, it follows that $\alpha\in S=-\alpha+S$, whence $0=\alpha-\alpha\in S$. The converse holds since $0\in{\mathcal{G}}(S)^{+}$ is impossible.
2): Assume that there is no $\gamma\in G$ such that ${\mathcal{G}}(S)^{+}=\gamma+S$. Take any $\alpha\in G$ such that $\alpha+{\mathcal{G}}(S)^{+}\subseteq S$, i.e., ${\mathcal{G}}(S)^{+}\subseteq-\alpha+S$. Then by assumption, ${\mathcal{G}}(S)^{+}\subsetneq-\alpha+S$. Since ${\mathcal{G}}(-\alpha+S)={\mathcal{G}}(S)$ by part 3) of Lemma 2.14, we have ${\mathcal{G}}(-\alpha+S)^{+}={\mathcal{G}}(S)^{+}$. Then $0\in-\alpha+S$ by part 1) of our Proposition, so $\alpha\in S$.
Now assume that ${\mathcal{G}}(S)^{+}=\gamma+S$ for some $\gamma\in G$. Then $-\gamma+{\mathcal{G}}(S)^{+}=S$, but $-\gamma\notin S$ since otherwise $0\in{\mathcal{G}}(S)^{+}$, which is impossible.
3): Take $\beta\in{\mathcal{G}}(S_{1})\setminus{\mathcal{G}}(S_{2})$ with $\beta>0$. Then $\beta+S_{2}\subsetneq S_{2}$ and there is $\alpha\in S_{2}\setminus\beta+S_{2}\,$. Now $S_{1}+S_{2}=S_{1}+\beta+S_{2}\subseteq S_{1}+\alpha^{-}\subseteq S_{1}+S_{2}\,$, whence $S_{1}+S_{2}=S_{1}+\alpha^{-}=S_{1}+\alpha$.
It is also possible to define the invariance group of a quasi-cut.
@@ 2.5. Partial answers to questions (QFS1), (QFS2) and (QFS3)
Since $S_{1}+T_{0}=S_{2}\,$, by part 4) of Lemma 2.6 also $S_{1}+\widehat{T_{0}}=S_{2}$ holds. Suppose that $T^{\prime}$ is a larger solution, and pick some $\tau^{\prime}\in T^{\prime}\setminus\widehat{T_{0}}$. If $\widehat{T_{0}}$ has a minimum, then denote it by $\tau$. If $\widehat{T_{0}}$ has no minimum, then $\widehat{T_{0}}=T_{0}$ has no infimum and $G\setminus T_{0}$ has no supremum, so there is $\tau\in G$ such that $\tau^{\prime}<\tau<T_{0}\,$. In both cases, we have that $\tau\leq\widehat{T_{0}}$. Consequently,
$S_{2}\>=\>S_{1}+\widehat{T}_{0}\>\subseteq\>S_{1}+\tau\>\subseteq\>S_{1}+\tau^{\prime}\>\subseteq\>S_{1}+T^{\prime}\>=\>S_{2}\>,$
so equality holds everywhere. It follows that $S_{2}+\tau^{\prime}-\tau=S_{1}+\tau+\tau^{\prime}-\tau=S_{1}+\tau^{\prime}=S_{2}\,$, whence $\tau^{\prime}-\tau\in{\mathcal{G}}(S_{2})=\{0\}$. Thus $\tau^{\prime}=\tau$, a contradiction. ∎
@@
Take final segments $S_{1}$ and $S_{2}$ of $G$. Assume that ${\mathcal{G}}(S_{1})={\mathcal{G}}(S_{2})=\{0\}$ and that $S_{2}$ is nonprincipal if $S_{1}$ is nonprincipal. Then
(22) $S_{1}\,-_{\rm fs}\,S_{2}\>=\>\left\{\begin{array}[]{ll}S_{2}-\alpha&\mbox{if $S_{1}=\alpha^{-}$ for some $\alpha\in G$, }\\ (S_{2}-S_{1}^{c})\,\widehat{}&\mbox{if $S_{1}$ is nonprincipal.}\end{array}\right.$
@@
As stated in the introduction, $T=S_{1}-_{\rm fs}S_{2}$ is the largest solution of $S_{1}+T\subseteq S_{2}$; hence if a solution of equation (1) exists, then $T=S_{1}-_{\rm fs}S_{2}$ is the largest solution of it. For $S_{1}=\alpha^{-}$, our assertion follows from Lemma 2.12. If $S_{1}$ is nonprincipal, then by assumption, $S_{2}$ is nonprincipal, and it follows from part 2) of Corollary 2.19 that $S_{2}-S_{1}^{c}$ is a solution of (1). Hence our statement for this case follows from Lemma 2.22. ∎

### v3

@@ 2.3. Subtraction of final segments
The final segment $S_{2}-_{\rm mc}S_{1}$ is closed, except in the following cases. If $S_{1}=\alpha^{-}$ and $S_{2}=\beta^{+}$ for some $\alpha,\beta\in G$ and $\beta$ has no immediate successor in $G$, then $\beta-\alpha$ has no immediate successor in $G$ and therefore, $S_{2}-_{\rm mc}S_{1}=(\beta-\alpha)^{+}$ is not closed. If $S_{2}$ and $S_{1}$ have no infimum in $G$, then $S_{2}-_{\rm mc}S_{1}$ may or may not have an infimum in $G$. For instance, if $G=\mathbb{Q}$ and $S_{1}=S_{2}=\{\alpha\in\mathbb{Q}\mid\alpha>\pi\}$, then $\Delta S_{1}=\{\alpha\in\mathbb{Q}\mid\alpha>-\pi\}$ and $S_{2}-_{\rm mc}S_{1}=0^{+}$ which is not closed although $S_{1}$ and $S_{2}$ are.
The question arises whether the difference operation $-_{\rm mc}$ provides some positive answer to our questions (QFS1) and (QFS2). As part 2) of Proposition 2.5 in conjunction with part 4) of Lemma 2.10 shows, it indeed gives a complete answer to these questions in the case where $S_{1}$ is principal. It now remains to deal with the case where $S_{1}$ is nonprincipal. This case is much more complex and we need some preparation.
@@ 2.4. Invariance groups
Take any non-empty subset $M\subseteq G$.
1)  We have that ${\mathcal{G}}(M)$ is a subgroup of $G$.
@@
4)  We have $0\in M$ if and only if ${\mathcal{G}}(M)\subseteq M$.
@@
(25) ${\mathcal{G}}(\alpha^{-})\>=\>{\mathcal{G}}(\alpha^{+})\>=\>{\mathcal{G}}(\{\alpha\})\>=\>\{0\}\>.$
@@
4):  If $0\in M$, then ${\mathcal{G}}(M)=0+{\mathcal{G}}(M)\subseteq M+{\mathcal{G}}(M)=M$. The converse is trivial.
@@
6):  Assume that $\alpha\in G$ is the infimum of $M$ in $G$. Then for every positive $\gamma\in G$ there is $\beta\in M$ such that $0\leq\beta-\alpha<\gamma$. Hence $\beta-\gamma<\alpha$ so that $\beta-\gamma\notin M$ and thus $\gamma\notin{\mathcal{G}}(M)$ by (24). This proves that ${\mathcal{G}}(M)=\{0\}$. ∎
@@
for every two convex subgroups $H_{1}$ and $H_{2}$ of $G$.
@@
${\mathcal{G}}(S_{1}+S_{2})\>=\>{\mathcal{G}}(S_{1})+{\mathcal{G}}(S_{2})\>=\>{\mathcal{G}}(S_{1})\cup{\mathcal{G}}(S_{2})\>.$
3)  If $\alpha\in G$ and $S$ is a final segment of $G$, then
@@
3):  We compute: ${\mathcal{G}}(\alpha+S)={\mathcal{G}}(\alpha^{-}+S)={\mathcal{G}}(\alpha^{-})\cup{\mathcal{G}}(S)=\{0\}\cup{\mathcal{G}}(S)={\mathcal{G}}(S)$, where we have used part 2) of our lemma together with equations (11) and (25)
@@
6):  Since ${\mathcal{G}}(S)$ is a convex subgroup of $G$, ${\mathcal{G}}(S)^{-}$ is the disjoint union of ${\mathcal{G}}(S)$ and ${\mathcal{G}}(S)^{+}$. If $\alpha\in{\mathcal{G}}(S)$, then $\alpha+S=S$. If $\alpha\in{\mathcal{G}}(S)^{+}$, then $\alpha>0$ and therefore, $\alpha+S\subseteq S$; but as $\alpha\notin{\mathcal{G}}(S)$, $\alpha+S\neq S$. This proves the implications $\Leftarrow$” in both statements.
@@
Note that the set $\{n\alpha\mid\alpha\in S\}$ is a final segment if and only if $G$ is $n$-divisible; in general, $nS$ is the smallest final segment that contains $\{n\alpha\mid\alpha\in S\}$. By a repeated application of part 2) of the last lemma, we obtain:
@@
From part 2) of Lemma 2.13 and the fact that the convex subgroups of $G$ are linearly ordered by inclusion, we obtain:
@@
Take final segments $S_{1}\,$ $S_{2}$ and $T$ such that $S_{1}+T=S_{2}$. Then ${\mathcal{G}}(S_{1})\subseteq{\mathcal{G}}(S_{1})\cup{\mathcal{G}}(T)={\mathcal{G}}(S_{2})$, and if ${\mathcal{G}}(S_{1})\subsetneq{\mathcal{G}}(S_{2})$, then ${\mathcal{G}}(T)={\mathcal{G}}(S_{2})$.
@@
Take a final segment $S$ of $G$.
@@
2) If ${\mathcal{G}}(S)\neq\{0\}$ then $S$ is nonprincipal.
@@
1): The first and second equality are special cases of part 3) of Lemma 2.12. The equality ${\mathcal{G}}(S)={\mathcal{G}}(\Delta S)$ is seen as follows. If $S$ is nonprincipal, then by part 3) of Lemma 2.10, $\Delta S=-S^{c}$ and the equality follows from what we have already proved. If $S$ is principal, say $S=\alpha^{-}$ for some $\alpha\in G$, then $\Delta S=(-\alpha)^{-}$ by (20), whence ${\mathcal{G}}(\Delta S)=\{0\}={\mathcal{G}}(S)$ by (25).
2): If ${\mathcal{G}}(S)\neq\{0\}$, then there is a negative element $\alpha\in{\mathcal{G}}(S)$ so that for every $\beta\in S\,$, $\beta>\alpha+\beta\in S\,$; hence $S$ has no smallest element. ∎
This lemma enables us to also define the invariance groups of cuts and quasi-cuts.
@@
In order to give answers to questions (QFS1), (QFS2) and (QFS3) in the case where ${\mathcal{G}}(S_{1})={\mathcal{G}}(S_{2})=\{0\}$ in Section 2.5 below, we need some more preparation. We set $G^{>0}:=\{\gamma\in G\mid\gamma>0\}=0^{+}$.
Lemma 2.18.
Proof.
If $\gamma=\alpha-\beta$ with $\alpha\in S$ and $\beta\in S^{c}$, then $\alpha-\gamma=\beta\notin S$, so $-\gamma\notin{\mathcal{G}}(S)$. Since ${\mathcal{G}}(S)$ is a subgroup of $G$, this implies that $\gamma\notin{\mathcal{G}}(S)$. Conversely, if $H$ is a subgroup of $G$ properly containing ${\mathcal{G}}(S)$, then there is $\gamma\in H^{>0}\setminus{\mathcal{G}}(S)$. Then $-\gamma\notin{\mathcal{G}}(S)$ and there is $\alpha\in S$ such that $\alpha-\gamma\notin S$. This implies $\beta=\alpha-\gamma\in S^{c}$, whence $\gamma=\alpha-\beta\in S-S^{c}$.
Now we prove equation (26). Since $\alpha>\beta$ whenever $\alpha\in S$ and $\beta\in S^{c}$, it follows that $S-S^{c}\subseteq G^{>0}\setminus{\mathcal{G}}(S)$. For the converse, assume that $\gamma\in G^{>0}\setminus{\mathcal{G}}(S)$. Then there is some $\beta\in S^{c}$ such that $\beta+\gamma\in S$ and therefore, $\gamma\in S-S^{c}$. ∎
Corollary 2.19.
1) We have
$S_{1}\,-_{\rm mc}\,S_{1}\>=\>S_{1}-S_{1}^{c}\>=\>{\mathcal{G}}(S_{1})^{+}\>,$
so that
(27) $S_{1}\,+\,(S_{2}-S_{1}^{c})\>=\>S_{1}\,+\,(S_{2}\,-_{\rm mc}\,S_{1})\>=\>{\mathcal{G}}(S_{1})^{+}+S_{2}\>\subseteq S_{2}\>.$
Proof.
1): This follows from equation (21) in part 3) of Lemma 2.10 together with Lemma 2.18. The equalities in (27) follow from part 1) together with parts 3) and 5) of Lemma 2.10. The inequality holds since ${\mathcal{G}}(S_{1})^{+}$ only contains positive elements.
2): This follows from part 1) together with part 3) of Lemma 2.6
3): Since ${\mathcal{G}}(S_{1})\subsetneq{\mathcal{G}}(S_{2})$, there is some $\alpha\in{\mathcal{G}}(S_{1})^{+}\cap{\mathcal{G}}(S_{2})$. We obtain that $S_{2}=\alpha+S_{2}\subseteq{\mathcal{G}}(S_{1})^{+}+S_{2}\,$. In view of (27), this proves statement 3). ∎
Finally, for application in Section 3.3 we prove the following facts:
Proposition 2.20.
Take final segments $S$, $S_{1}$ and $S_{2}$ of $G$.
1) We have that ${\mathcal{G}}(S)^{+}$ does not contain $S$ if and only if $0\in S$.
Proof.
1): If ${\mathcal{G}}(S)^{+}$ does not contain $S$, then $S\setminus{\mathcal{G}}(S)^{+}\neq\emptyset$, so there is some $\alpha\in S\cap{\mathcal{G}}(S)$. Since ${\mathcal{G}}(S)$ is a group, it follows that $-\alpha\in{\mathcal{G}}(S)$. Therefore $S=-\alpha+S$, whence $0=\alpha-\alpha\in S$. The converse holds since $0\in{\mathcal{G}}(S)^{+}$ is impossible.
2): Since ${\mathcal{G}}(S)^{+}$ only contains positive elements, $S$ is always contained in $\{\alpha\in G\mid\alpha+{\mathcal{G}}(S)^{+}\subseteq S\}$. Assume that there is no $\gamma\in G$ such that ${\mathcal{G}}(S)^{+}=\gamma+S$. Take any $\alpha\in G$ such that $\alpha+{\mathcal{G}}(S)^{+}\subseteq S$, i.e., ${\mathcal{G}}(S)^{+}\subseteq-\alpha+S$. Then by assumption, ${\mathcal{G}}(S)^{+}\subsetneq-\alpha+S$. Since ${\mathcal{G}}(-\alpha+S)={\mathcal{G}}(S)$ by part 3) of Lemma 2.13, we have ${\mathcal{G}}(-\alpha+S)^{+}={\mathcal{G}}(S)^{+}$. Then $0\in-\alpha+S$ by part 1) of our Proposition, so $\alpha\in S$.
Now assume that ${\mathcal{G}}(S)^{+}=\gamma+S$ for some $\gamma\in G$. Then $-\gamma+{\mathcal{G}}(S)^{+}\subseteq S$, but $-\gamma\notin S$ since otherwise $0\in{\mathcal{G}}(S)^{+}$, which is impossible.
3): Take $\beta\in{\mathcal{G}}(S_{1})\setminus{\mathcal{G}}(S_{2})$ with $\beta>0$. Then $\beta+S_{2}\subsetneq S_{2}$ and there is $\alpha\in S_{2}\setminus\beta+S_{2}\,$, so that $\beta+S_{2}\subseteq\alpha^{-}\subseteq S_{2}\,$. Now $S_{1}+S_{2}=(S_{1}+\beta)+S_{2}=S_{1}+(\beta+S_{2})\subseteq S_{1}+\alpha^{-}\subseteq S_{1}+S_{2}\,$, whence $S_{1}+S_{2}=S_{1}+\alpha^{-}=S_{1}+\alpha$.
@@ 2.5. Partial answers to questions (QFS1), (QFS2) and (QFS3)
Since $S_{1}+T_{0}=S_{2}\,$, by part 4) of Lemma 2.6 also $S_{1}+\widehat{T_{0}}=S_{2}$ holds. Suppose that $T^{\prime}$ is a larger solution, and pick some $\alpha^{\prime}\in T^{\prime}\setminus\widehat{T_{0}}$. If $\widehat{T_{0}}$ has a minimum, then denote it by $\alpha$. If $\widehat{T_{0}}$ has no minimum, then $\widehat{T_{0}}=T_{0}$ has no infimum and $G\setminus T_{0}$ has no supremum, so there is $\alpha\in G$ such that $\alpha^{\prime}<\alpha<T_{0}\,$. In both cases, we have that $\alpha\leq\widehat{T_{0}}$. Consequently,
$S_{2}\>=\>S_{1}+\widehat{T}_{0}\>\subseteq\>S_{1}+\alpha\>\subseteq\>S_{1}+\alpha^{\prime}\>\subseteq\>S_{1}+T^{\prime}\>=\>S_{2}\>,$
so equality holds everywhere. It follows that $S_{2}+\alpha^{\prime}-\alpha=S_{1}+\alpha+\alpha^{\prime}-\alpha=S_{1}+\alpha^{\prime}=S_{2}\,$, whence $\alpha^{\prime}-\alpha\in{\mathcal{G}}(S_{2})=\{0\}$. Thus $\alpha^{\prime}=\tau$, a contradiction. ∎
@@
Take final segments $S_{1}$ and $S_{2}$ of $G$. Assume that ${\mathcal{G}}(S_{1})={\mathcal{G}}(S_{2})=\{0\}$ and that $S_{2}$ is nonprincipal if $S_{1}$ is nonprincipal. Then $T=S_{2}-_{\rm fs}S_{1}$ is the largest solution of (1), and we have
(28) $S_{2}\,-_{\rm fs}\,S_{1}\>=\>\left\{\begin{array}[]{ll}S_{2}-\alpha&\mbox{if $S_{1}=\alpha^{-}$ for some $\alpha\in G$, }\\ (S_{2}-S_{1}^{c})\,\widehat{}&\mbox{if $S_{1}$ is nonprincipal.}\end{array}\right.$
@@
As stated in the introduction, $T=S_{2}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T\subseteq S_{2}$; hence if a solution of equation (1) exists, then $T=S_{2}-_{\rm fs}S_{1}$ is the largest solution of it. For $S_{1}=\alpha^{-}$, our assertion follows from Lemma 2.5. If $S_{1}$ is nonprincipal, then by assumption, $S_{2}$ is nonprincipal, and it follows from part 2) of Corollary 2.19 that $S_{2}-S_{1}^{c}$ is a solution of (1). Hence our statement for this case follows from Lemma 2.21. ∎